# On an entropy inequality for quadratic forms and applications

### v1

@@ 2.4 Frostman measures
Let $\mu_{X}$, and $\mu_{Y}$ be the distribution measures associated to $X$ and $Y$, respectively.
(i) When $b\neq 0$, for any ball $B(y^{\prime},r)$ one has
[…]
A similar conclusion follows when $a\neq 0$. When $a=b=0$, the conclusion follows automatically.
(ii) For any ball $B((x,y),r)$,
$P((aX+bY,cX+dY)\in B((x,y),r)$ $\leq P(aX+bY\in B(x,r),cX+dY\in B(y,r)).$
If both $aX+bY\in B(x,r)$ and $cX+dY\in B(y,r)$ occur, we must have $X\in B\left(\frac{dx-by}{ad-bc},\frac{(|d|+|b|)r}{|ad-bc|}\right)$ and $Y\in B\left(\frac{-cx+ay}{ad-bc},\frac{(|c|+|a|)r}{|ad-bc|}\right)$. Therefore,
$P(aX+bY\in B(x,r),cX+dY\in B(y,r))\leq C\left(\frac{(|d|+|b|)r}{|ad-bc|}\right)^{s}\cdot C\left(\frac{(|c|+|a|)r}{|ad-bc|}\right)^{s}=O_{a,b,c,d}(C^{2}r^{2s}).$
The lemma now follows. ∎
The next property is taken from [11, Lemma 4.8].
@@
If $X$ is $(s,C)$-Frostman for some $s,C>0$, then $H_{n}(X)\geq ns-\log C+O(1)$.
@@
Let $Z$ be a real random variable with bounded range. Assume that there is some $s>0$ and $C\geq 1$ such that $P(|Z|<r)\leq Cr^{s}$ for any $r>0$. For any $0<\eta\leq 2^{-1}(2C)^{-1/s}$, denote by $E$ the event that $|Z|\geq\eta$. Then the random variable $Z_{E}$ satisfies
$H_{n}(Z_{E})=\frac{H_{n}(Z)}{P(E)}+O(C(2\eta)^{s}n)+O(1).$
@@
Recall that $Z_{E}$ is the random variable such that $Z_{E}\equiv(Z|E)$, and $P_{E}$ is the conditional probability measure given $E$.
Since $P(|Z|<\eta)\leq C\eta^{s}\leq 1/2$, we have $P(E)\geq 1-C\eta^{s}\geq 1/2$. Assume that $k_{0}$ is a natural number with $(k_{0}-1)2^{-n}\leq\eta<k_{0}2^{-n}$. Combining the conditions in this lemma, one deduces that
$H_{n}(Z_{E})$ $=-\sum\limits_{|k|\geq k_{0}+1}P_{E}(D_{n}(Z_{E})=k2^{-n})\log P_{E}(D_{n}(Z_{E})=k2^{-n})$
$\qquad+O\left(-\sum\limits_{k_{0}-1\leq|k|\leq k_{0}}P_{E}(D_{n}(Z_{E})=k2^{-n})\log P_{E}(D_{n}(Z_{E})=k2^{-n})\right)$
$=-\sum\limits_{|k|\geq k_{0}+1}\frac{P(D_{n}(Z)=k2^{-n})}{P(E)}\big{(}\log P(D_{n}(Z)=k2^{-n})-\log P(E)\big{)}+O(1)$
$=-\frac{1}{P(E)}\sum\limits_{|k|\geq k_{0}+1}P(D_{n}(Z)=k2^{-n})\log P(D_{n}(Z)=k2^{-n})+O(1).$
First, let us consider the case that $\eta>2^{-n+1}$, which gives $(k_{0}+1)2^{-n}\leq\eta+2^{-n+1}<2\eta$. Applying Lemma 2.10 with the fact
$\sum\nolimits_{|k|\leq k_{0}}P(D_{n}(Z)=k2^{-n})\leq P(|Z|<(k_{0}+1)2^{-n})\leq P(|Z|<2\eta)\leq C(2\eta)^{s}\leq 1/2,$
@@
$-\sum\limits_{|k|\leq k_{0}}P(D_{n}(Z)=k2^{-n})\log P(D_{n}(Z)=k2^{-n})$
$\quad\quad\quad\leq C(2\eta)^{s}\cdot\log\big{(}2\eta 2^{n}+3\big{)}+O(1)\leq C(2\eta)^{s}n+O(1).$
Second, we consider the case that $\eta\leq 2^{-n+1}$. Then $k_{0}\leq 3$ and
$-\sum\nolimits_{|k|\leq k_{0}}P(D_{n}(Z)=k2^{-n})\log P(D_{n}(Z)=k2^{-n})=O(1).$
@@
$H(Z_{E})$ $=\frac{1}{P(E)}\left(H_{n}(Z)+\sum\limits_{|k|\leq k_{0}}P(D_{n}(Z)=k2^{-n})\log P(D_{n}(Z)=k2^{-n})\right)+O(1)$
$=\frac{H_{n}(Z)}{P(E)}+O(C(2\eta)^{s}n)+O(1).$
@@
Let $Z$, $Z^{\prime}$ be independent random variables. Suppose that $Z$ is $(s,C)$-Frostman, where $s,C>0$. Then
@@
[…]
@@
[…]
for any Borel set $B$. That is to say, $\Delta(\mu)$ is the image of $\mu\times\mu$ under the distance map $(x,y)\mapsto|x-y|$. We now recall a result from Mattila in [12, Section 4.2, page 61].
@@
Let $\mu$ be a $(2s,C)$-Frostman probability measure with $\operatorname{spt}(\mu)\subset B(0,R)\subset\mathbb{R}^{2}$, where $\frac{1}{4}<s\leq 1$, $C\geq 1$ and $R>0$. Then for any $r,\eta,t$ with $0<\eta<r<2R$, $0<t\leq 1$ and $t<2s-\frac{1}{2}$, it satisfies that
$\Delta(\mu)([r,r+\eta])=O_{t}\left(\frac{CR^{2s-t}\eta^{t}}{2s-t-1/2}\right)+O_{t}(R^{-t}\eta^{t}).$
@@
Since $\mu$ is $(2s,C)$-Frostman, for $20<t<2s-1/$ 2 and any $x\in\mathbb{R}^{2}$,
[…]
$\qquad=O_{t}\left(\frac{CR^{2s-t-1/2}}{2s-t-1/2}\right)+O_{t}(R^{-t-1/2}):=M.$
@@
$\Delta(\mu)([r,r+\eta])\ll_{t}r^{1/2}\eta^{t}I_{t+1/2}(\mu)=O_{t}\left(\frac{CR^{2s-t}\eta^{t}}{2s-t-1/2}\right)+O_{t}(R^{-t}\eta^{t}).$
2.5 Entropy of the distance function
@@
Let $W$ be a $(2s,C)$-Frostman random variable with $range(W)\subseteq B(0,R)\subset\mathbb{R}^{2}$, where $11/2<s<$ 1, $C\geq 1$ and $R>0$. Let $W^{\prime}$ be a random variable which is i.i.d with $W$. Then for any number $t$ satisfying $t\leq 1$ and $s<t<2s-1/2$, we have
@@
$p_{k}$ $=\Delta(\mu)\big{(}[k2^{-n},(k+1)2^{-n}\big{)}=O_{t}\left(\frac{CR^{2s-t}2^{-nt}}{2s-t-1/2}\right)+O_{t}(R^{-t}2^{-nt}).$
@@
The parameters $C$ and $R$ in Theorem 2.32 are allowed to be unbounded as $n$ tends to infinity. In Example A.5, we show an example in which $C$ is growing with $n$. This example will tell us that the term $-\log C$ on the right-hand side of (14) is essential.
@@
When $13/4<s<$ 1, we can take $t=1$ in Theorem 2.32. Combining this with (11), we then obtain the identity
$H_{n}(|W-W^{\prime}|)=n+O(\log C)+O(|\log R|)+O_{s}(1).$
@@
Let $W$ be a $(2s,C)$-Frostman random variable with $range(W)\subseteq B(0,R)\subset\mathbb{R}^{2}$, where $11/2<s<$ 1, $C\geq 1$ and $R>0$. Let $W^{\prime}$ be a random variable which is i.i.d with $W$. Then there exists a constant $\epsilon_{0}=\epsilon_{0}(s)>0$ such that
$H_{n}(|W-W^{\prime}|^{2})\geq n(s+\epsilon_{0})+O(\log C)+O(|\log R|)+O_{s}(1)$
@@
$H_{n}(Z)\geq nt-\log C+O(|\log R|)+O(|\log(2s-t-1/2)|)+O_{t}(1).$
By Lemma 2.28, one has $P(|Z|<r)\leq Cr^{2s}$ for any $r>0$. Let $\eta$ be a parameter with $0<\eta\leq 2^{-1}(2C)^{-1/2s}$ to be determined later. Denote by $E$ the event that $|Z|\geq\eta$. By Lemma 2.27, we deduce that
@@
Applying Lemma B.2, one deduces that
@@
$\geq P(E)\left(\frac{nt}{P(E)}+O(\log C)+O(|\log R|)\right.$
@@
Now we take $t=s+\frac{2s-1}{3}$ when $41/2<s\leq 3/$ 4, and take $t=1$ when $13/4<s<$ 1. Moreover, assume that the implied constant in the term $O(C\eta^{2s}n)$ in (15) is $c_{0}$, with $c_{0}\geq 2$. We take $\eta$ to be the number satisfying
$c_{0}C\eta^{2s}=\begin{cases}(2s-1)/6,\quad&\text{if }1/2<s\leq 3/4,\\ (2-2s)/3,\quad&\text{if }3/4<s<1.\\ \end{cases}$
One can verify that the requirements involving $t$ and $\eta$ are satisfied, and $12s-t-1/2\gg_{s}$ 1. Also, $\log(1/\eta)<\log C+O_{s}(1)$. As a result, we conclude that
$H_{n}(Z^{2})\geq n(s+\epsilon_{0})+O(\log C)+O(|\log R|)+O_{s}(1),$
with $\epsilon_{0}=(2s-1)/6$ when $41/2<s\leq 3/$ 4, and $\epsilon_{0}=(1-s)/3$ when $13/4<s<$ 1. ∎
@@
Let $X$, $X^{\prime}$, $Y$, $Y^{\prime}$ be independent real random variables. Suppose that $X$ and $X^{\prime}$ are both $(s,C)$-Frostman with ranges contained in $[-c,c]$, where $0<s\leq 1$ and $C,c\geq 1$. Then, for any $n\in\mathbb{N}$,
$H_{n}\big{(}(X-X^{\prime})^{2}+(Y-Y^{\prime})^{2}\big{)}\geq ns+O_{s}(\log C)+O_{s}(\log n),$
@@
$H_{n}\big{(}\sqrt{(X-X^{\prime})^{2}+(Y-Y^{\prime})^{2}}\big{)}\geq ns+O_{s}(\log C)+O(\log c)+O_{s}(\log n).$
@@
$H(Z^{2})\geq ns+O(\log C)+O(C\eta^{s}n)+O\big{(}\log\frac{1}{\eta}\big{)}+O(1).$
The first inequality follows by taking $\eta=2^{-1}\big{(}\frac{\log n}{2Cn}\big{)}^{1/s}$. The second inequality follows from the fact that
$H_{n}\big{(}(X-X^{\prime})^{2}+(Y-Y^{\prime})^{2}\big{)}\leq H_{n}\big{(}\sqrt{(X-X^{\prime})^{2}+(Y-Y^{\prime})^{2}}\big{)}+O(\log c).$

### v2

@@ 2.4 Frostman measures
By Lemma 2.22, the random variable $(X,Y)$ is marginally $(s_{1},C_{1};s_{2},C_{2})$-Frostman.
(i) Consider any ball $B(z,r)$. Let $\{I_{i}\}$ be a partition of $\mathbb{R}$} composed of intervals of length $r$. Then
$P(aX+bY\in B(z,r))=\sum\limits_{i}P(aX+bY\in B(z,r),\,X\in I_{i}):=\sum\limits_{i}q_{i}.$
Assume that $b\neq 0$. If $q_{i}>0$, then $Y$ is located in an interval $J_{i}$ of length at most $\frac{2(|a|+1)r}{|b|}$. So
$q_{i}\leq P(X\in I_{i},\,Y\in J_{i})\leq C_{2}\Big{(}\frac{(|a|+2)r}{|b|}\Big{)}^{s_{2}}P(X\in I_{i}).$
It follows that
$P(aX+bY\in B(z,r))\leq\sum\limits_{i}C_{2}\Big{(}\frac{(|a|+2)r}{|b|}\Big{)}^{s_{2}}P(X\in I_{i})=C_{2}\Big{(}\frac{(|a|+2)r}{|b|}\Big{)}^{s_{2}}.$
Hence $aX+bY$ is $(s_{2},C^{\prime}_{2})$-Frostman with $\log C^{\prime}_{2}=\log C_{2}+O_{a,b}(1)$. By symmetry, it is also $(s_{1},C^{\prime}_{1})$-Frostman with $\log C_{1}^{\prime}=\log C_{1}+O_{a,b}(1)$, provided that $a\neq 0$.
(ii) Write $p:=P\big{(}aX+bY\in B(z_{1},r_{1}),\,cX+dY\in B(z_{2},r_{2})\big{)}$ for an arbitrary pair of balls $B(z_{1},r_{1})$ and $B(z_{2},r_{2})$. Calculation shows that, if $p>0$, then we must have
$X\in B\Big{(}\frac{dz_{1}-bz_{2}}{ad-bc},\,\frac{|d|r_{1}+|b|r_{2}}{|ad-bc|}\Big{)},\qquad Y\in B\Big{(}\frac{-cz_{1}+az_{2}}{ad-bc},\,\frac{|c|r_{1}+|a|r_{2}}{|ad-bc|}\Big{)}.$ (17)
First, let us consider the case $ad\neq 0$. When $r_{1}\leq r_{2}$, let $\{J_{i}\}$ be a partition of $\mathbb{R}$} composed of intervals of length $r_{1}$. We estimate $p$ by
$p$ $=\sum\limits_{i}P\big{(}aX+bY\in B(z_{1},r_{1}),\,cX+dY\in B(z_{2},r_{2}),\,Y\in J_{i}\big{)}$
$\leq\sum\limits_{i}P\big{(}aX+bY\in B(z_{1},r_{1}),\,Y\in J_{i}\big{)}:=\sum\limits_{i}p_{i}.$
One can check that, if $p_{i}>0$, then $X$ is located in an interval $I_{i}$ of length at most $\frac{2(|b|+1)r_{1}}{|a|}$. Since $(X,Y)$ is conditionally $(s_{1},C_{1};s_{2},C_{2})$-Frostman, we have
$p_{i}\leq P(X\in I_{i},Y\in J_{i})\leq C_{1}\Big{(}\frac{(|b|+2)r_{1}}{|a|}\Big{)}^{s_{1}}P(Y\in J_{i}).$
Combining (17), it follows that
$p$ $\leq\sum\limits_{i}C_{1}\Big{(}\frac{(|b|+2)r_{1}}{|a|}\Big{)}^{s_{1}}P(Y\in J_{i})$
$\leq C_{1}\Big{(}\frac{(|b|+2)r_{1}}{|a|}\Big{)}^{s_{1}}P\left(Y\in B\Big{(}\frac{-cz_{1}+az_{2}}{ad-bc},\,\frac{(|c|+|a|)r_{2}}{|ad-bc|}\Big{)}\right)$
$\leq C_{1}\Big{(}\frac{(|b|+2)r_{1}}{|a|}\Big{)}^{s_{1}}\cdot C_{2}\Big{(}\frac{(|c|+|a|)r_{2}}{|ad-bc|}\Big{)}^{s_{2}}.$
When $r_{1}>r_{2}$, let $\{I_{i}^{\prime}\}$ be a partition of $\mathbb{R}$} composed of intervals of length $r_{2}$. We estimate $p$ by
$p\leq\sum\limits_{i}P\big{(}cX+dY\in B(z_{2},r_{2}),\,X\in I_{i}^{\prime}\big{)}:=\sum\limits_{i}p_{i}^{\prime}.$
We apply similar arguments as above. For non-zero terms, the value of $Y$ locates in an interval $J_{i}^{\prime}$ of length at most $\frac{2(|c|+1)r_{2}}{|d|}$. So
$p_{i}^{\prime}\leq P(X_{i}\in I_{i}^{\prime},Y\in J_{i}^{\prime})\leq C_{1}\Big{(}\frac{(|c|+2)r_{2}}{|d|}\Big{)}^{s_{2}}P(X\in I_{i}^{\prime}),$
and then
$p$ $\leq C_{1}\Big{(}\frac{(|c|+2)r_{2}}{|d|}\Big{)}^{s_{2}}P\left(X\in B\Big{(}\frac{dz_{1}-bz_{2}}{ad-bc},\,\frac{(|d|+|b|)r_{1}}{|ad-bc|}\Big{)}\right)$
$\leq C_{1}\Big{(}\frac{(|c|+2)r_{2}}{|d|}\Big{)}^{s_{2}}\cdot C_{1}\Big{(}\frac{dz_{1}-bz_{2}}{ad-bc},\,\frac{(|d|+|b|)r_{1}}{|ad-bc|}\Big{)}^{s_{1}}.$
We conclude that $(X,Y)$ is jointly $(s_{1},s_{2},C^{\prime\prime})$-Frostman with $\log C^{\prime\prime}=\log C_{1}C_{2}+O_{a,b,c,d}(1)$.
Second, when $ad=0$, then $bc\neq 0$ since $ad-bc\neq 0$. Similar arguments show that $(X,Y)$ is jointly $(s_{2},s_{1},C^{\prime\prime})$-Frostman. The proof is completed. ∎
The following lemma shows relations of the discretized entropies and the Frostman conditions.
@@
Let $s,s_{1},s_{2}\geq 0$, $C,C_{1},C_{2}>0$ and $X,Y,Z$ be random variables. The following statements hold for all $n\in\mathbb{N}$.
(i) If $Z$ is $(s,C)$-Frostman, then
$H_{n}(Z)\geq ns-\log C+O(1).$
(ii) If $X$ is conditionally $(s,C)$-Frostman given $Y$, then
$H_{n}(X|Y)\geq ns-\log C+O(1).$
Moreover, if $(X,Y)$ is conditionally $(s_{1},C_{1};s_{2},C_{2})$-Frostman, then
$H_{n}(X,Y)\geq\frac{1}{2}\big{(}H_{n}(X)+H_{n}(Y)+n(s_{1}+s_{2})-\log C_{1}C_{2}\big{)}+O(1).$
Proof.
(i) is taken from [12, Lemma 4.8].
(ii) Since $X$ is conditionally $(s,C)$-Frostman given $Y$, then
$H_{n}(X,Y)$ $=-\sum\limits_{x,y}P(D_{n}(X)=x,\,D_{n}(Y)=y)\log P(D_{n}(X)=x,\,D_{n}(Y)=y)$
$\geq-\sum\limits_{x,y}P(D_{n}(X)=x,\,D_{n}(Y)=y)\log\Big{(}O(C2^{-ns})\cdot P(D_{n}(Y)=y)\Big{)}$
$=\sum\limits_{x,y}P(D_{n}(X)=x,\,D_{n}(Y)=y)\cdot(ns-\log C+O(1))$
$\qquad\qquad-\sum\limits_{y}\big{(}\sum\limits_{x}P(D_{n}(X)=x,\,D_{n}(Y)=y)\big{)}\log P(D_{n}(Y)=y)$
$=ns-\log C+O(1)+H_{n}(Y),$
i.e., $H_{n}(X|Y)\geq ns-\log C+O(1)$.
For the conditionally $(s_{1},C_{1};s_{2},C_{2})$-Frostman random variable $(X,Y)$, we have both
$H_{n}(X,Y)-H_{n}(Y)\geq ns_{1}-\log C_{1}+O(1),\quad H_{n}(X,Y)-H_{n}(X)\geq ns_{2}-\log C_{2}+O(1).$
The conclusion then follows.
@@
Let $m\geq 1$. Suppose that $Z$ is a random variable with bounded range in ${\mathbb{R}}^{m}$, and $f$ is a function such that $domain(f)\supseteq range_{+}(Z)$. Assume that there are numbers $s,C>0$ such that $P(|f(Z)|<r)\leq Cr^{s}$ for any $r>0$. Let $0<\eta\leq\min\{2^{-1}(2C)^{-1/s},\,1/2\}$, and denote by $E$ the event that $|Z|\geq\eta$. Assume that $P(E)\gg 1$. Then the random variable $f(Z)_{E}$ satisfies
$H_{n}(f(Z)_{E})=\frac{H_{n}(f(Z))}{P(E)}+O(C(2\eta)^{s}n)+O_{m}(1).$
@@
Recall that $f(Z)_{E}$ is the random variable such that $f(Z)_{E}\equiv(f(Z)|E)$, and $P_{E}$ is the conditional probability measure given $E$ (see Lemma A.2).
Assume that $k_{0}$ is a natural number with $(k_{0}-1)2^{-n}\leq\eta<k_{0}2^{-n}$. Combining the conditions in this lemma, one deduces that
$H_{n}(f(Z)_{E})$ $=-\sum\limits_{k\in\mathbb{Z}^{m}\atop|k|\geq k_{0}+m}P_{E}(D_{n}(f(Z)_{E})=k2^{-n})\log P_{E}(D_{n}(f(Z)_{E})=k2^{-n})$
$\qquad+O\left(-\sum\limits_{k_{0}-m\leq|k|<k_{0}+m}P_{E}(D_{n}(f(Z)_{E})=k2^{-n})\log P_{E}(D_{n}(f(Z)_{E})=k2^{-n})\right)$
$=-\sum\limits_{|k|\geq k_{0}+m}\frac{P(D_{n}(f(Z))=k2^{-n})}{P(E)}\big{(}\log P(D_{n}(f(Z))=k2^{-n})-\log P(E)\big{)}+O_{m}(1)$
$=-\frac{1}{P(E)}\sum\limits_{|k|\geq k_{0}+m}P(D_{n}(f(Z))=k2^{-n})\log P(D_{n}(f(Z))=k2^{-n})+O_{m}(1).$
First, let us consider the case that $\eta>m2^{-n+2}$, which gives $(k_{0}+2m)2^{-n}\leq\eta+(2m+1)2^{-n}<2\eta$. Applying Lemma 2.9 with the fact
$\sum\nolimits_{|k|<k_{0}+m}P(D_{n}(f(Z))=k2^{-n})\leq P(|f(Z)|<(k_{0}+2m)2^{-n})\leq P(|f(Z)|<2\eta)\leq C(2\eta)^{s}\leq 1/2,$
@@
$-\sum\limits_{|k|<k_{0}+m}P(D_{n}(f(Z))=k2^{-n})\log P(D_{n}(f(Z))=k2^{-n})$
$\quad\quad\quad\leq C(2\eta)^{s}\cdot\log\big{(}2\eta 2^{n}+2m+3\big{)}+O(1)\leq C(2\eta)^{s}n+O_{m}(1).$
Second, we consider the case that $\eta\leq m2^{-n+2}$. Then $k_{0}\leq 4m+1$ and
$\sum\nolimits_{|k|<k_{0}+m}P(D_{n}(f(Z))=k2^{-n})\log P(D_{n}(f(Z))=k2^{-n})=O_{m}(1).$
@@
$H(f(Z)_{E})$ $=\frac{1}{P(E)}\left(H_{n}(f(Z))+\sum\limits_{|k|<k_{0}+m}P(D_{n}(f(Z))=k2^{-n})\log P(D_{n}(f(Z))=k2^{-n})\right)+O_{m}(1)$
$=\frac{H_{n}(f(Z))}{P(E)}+O(C(2\eta)^{s}n)+O_{m}(1).$
@@
Let $Z$, $Z^{\prime}$ be independent random variables. Suppose that $Z$ is $(s,C)$-Frostman, where $s\geq 0$ and $C>0$. Then
@@
$P(|Z-Z^{\prime}|<r)=\int P(|Z-z^{\prime}|<r)d\mu(z^{\prime})\leq\int Cr^{s}d\mu(z^{\prime})=Cr^{s}.$
2.5 Entropy of the distance function
@@
$\Delta(\mu)(B)=\int\mu(\{y\in\Omega:|x-y|\in B\})d\mu(x)$
for any Borel set $B$. That is to say, $\Delta(\mu)$ is the image of $\mu\times\mu$ under the distance map $(x,y)\mapsto|x-y|$. We now recall a result from Mattila in [13, Section 4.2, page 61].
@@
Let $\mu$ be a $(2s,C)$-Frostman probability measure with $\operatorname{spt}(\mu)\subset B(0,R)\subset\mathbb{R}^{2}$, where $\frac{1}{4}<s\leq 1$ and $C,R>0$. Then for any $r,\eta,t$ with $0<\eta<r<2R$, $0<t\leq 1$ and $t<2s-\frac{1}{2}$, it satisfies that
$\Delta(\mu)([r,r+\eta])=O_{t}\left(\Big{(}\frac{C\cdot\max\{R^{2s-t},1\}}{2s-t-1/2}+1\Big{)}\eta^{t}\right).$
@@
Put $\tilde{R}=\max\{R,1\}$. Since $\mu$ is $(2s,C)$-Frostman, for 0<t<2s1/20<t<2s-1/2 and any $x\in\mathbb{R}^{2}$,
$\int|x-y|^{-t-1/2}d\mu(y)=(t+1/2)\int_{0}^{\infty}\frac{\mu(B(x,\ell))}{\ell^{t+3/2}}d\ell$
$\qquad\leq(t+1/2)\left(\int_{0}^{2\tilde{R}}C\ell^{2s-t-3/2}d\ell+\int_{2\tilde{R}}^{\infty}\ell^{-t-3/2}d\ell\right)$
$\qquad=O_{t}\left(\frac{C\tilde{R}^{2s-t-1/2}}{2s-t-1/2}\right)+O_{t}(1):=M.$
@@
$\Delta(\mu)([r,r+\eta])\ll_{t}r^{1/2}\eta^{t}I_{t+1/2}(\mu)=O_{t}\left(\frac{C\tilde{R}^{2s-t}\eta^{t}}{2s-t-1/2}\right)+O_{t}(\eta^{t}).$
The proof is completed. ∎
@@
Let $W$ be a $(2s,C)$-Frostman random variable with $range(W)\subseteq B(0,R)\subset\mathbb{R}^{2}$, where $1/2<s<1$ and $C,R>0$. Let $W^{\prime}$ be a random variable which is i.i.d with $W$. Then for any number $t$ satisfying $t\leq 1$ and s<t<2s1/2s<t<2s-1/2, we have
$H_{n}(|W-W^{\prime}|)\geq nt-\log_{+}C+O(\log_{+}R)+O(|\log(2s-t-1/2)|)+O_{t}(1).$ (18)
@@
$p_{k}$ $=\Delta(\mu)\big{(}[k2^{-n},(k+1)2^{-n}\big{)}=O_{t}\left(\Big{(}\frac{C\cdot\max\{R^{2s-t},1\}}{2s-t-1/2}+1\Big{)}2^{-nt}\right).$
@@
The parameters $C$ and $R$ in Theorem 2.30 are allowed to be unbounded as $n$ tends to infinity. In Example 7.5, we show an example in which $C$ is growing with $n$. This example will tell us that the term $\log C$ on the right-hand side of (18) is essential.
@@
When $3/4<s<1$, we can take $t=1$ in Theorem 2.30. Combining this with (12), we then obtain the identity
$H_{n}(|W-W^{\prime}|)=n+O(\log_{+}C)+O(\log_{+}R)+O_{s}(1).$
@@
Let $W$ be a $(2s,C)$-Frostman random variable with $range(W)\subseteq B(0,R)\subset\mathbb{R}^{2}$, where $1/2<s<1$ and $C,R>0$. Let $W^{\prime}$ be a random variable which is i.i.d with $W$. Then there exists a constant $\epsilon_{0}=\epsilon_{0}(s)>0$ such that
$H_{n}(|W-W^{\prime}|^{2})\geq n(s+\epsilon_{0})+O(\log_{+}C)+O(\log_{+}R)+O_{s}(1)$
@@
$H_{n}(Z)\geq nt+O(\log_{+}C)+O(\log_{+}R)+O(|\log(2s-t-1/2)|)+O_{t}(1).$
By Lemma 2.26, one has $P(|Z|<r)\leq Cr^{2s}$ for any $r>0$. Let $\eta$ be a parameter with $0<\eta<\min\{2^{-1}(2C)^{-1/2s},\,1/2\}$ to be determined later. Denote by $E$ the event that $|Z|\geq\eta$. By Lemma 2.25, we deduce that
@@
Applying Lemma A.2, one deduces that
@@
$\geq P(E)\left(\frac{nt}{P(E)}+O(\log_{+}C)+O(\log_{+}R)\right$.
@@
Now we take $t=s+\frac{2s-1}{3}$ when $1/2<s\leq 3/4$, and take $t=1$ when $3/4<s<1$. Moreover, assume that the implied constant in the term $O(C\eta^{2s}n)$ in (19) is $c_{0}$, with $c_{0}\geq 2$. We take $\eta$ to be the number satisfying
[…]
One can verify that the requirements involving $t$ and $\eta$ are satisfied, and $2s-t-1/2\gg_{s}1$. Also, $\log(1/\eta)<\log C+O_{s}(1)$. As a result, we conclude that
$H_{n}(Z^{2})\geq n(s+\epsilon_{0})+O(\log_{+}C)+O(\log_{+}R)+O_{s}(1),$
with $\epsilon_{0}=(2s-1)/6$ when $1/2<s\leq 3/4$, and $\epsilon_{0}=(1-s)/3$ when $3/4<s<1$. ∎
@@
Let $X$, $X^{\prime}$, $Y$, $Y^{\prime}$ be independent real random variables. Suppose that $X$ and $X^{\prime}$ are both $(s,C)$-Frostman with ranges contained in [c,c][-c,c], where $0<s\leq 1$ and $C,c>0$. Then, for any $n\in\mathbb{N}$,
$H_{n}\big{(}(X-X^{\prime})^{2}+(Y-Y^{\prime})^{2}\big{)}\geq ns+O_{s}(\log_{+}C)+O_{s}(\log n),$
@@
$H_{n}\big{(}\sqrt{(X-X^{\prime})^{2}+(Y-Y^{\prime})^{2}}\big{)}\geq ns+O_{s}(\log_{+}C)+O(\log_{+}c)+O_{s}(\log n).$
@@
$H(Z^{2})\geq ns+O(\log_{+}C)+O(C\eta^{s}n)+O\big{(}\log\frac{1}{\eta}\big{)}+O(1).$
The first inequality follows by taking $\eta=\min\left\{\frac{1}{2}\big{(}\frac{\log n}{2Cn}\big{)}^{1/s},\,\frac{1}{10}\right\}$. The second inequality follows from the fact that
$H_{n}\big{(}(X-X^{\prime})^{2}+(Y-Y^{\prime})^{2}\big{)}\leq H_{n}\big{(}\sqrt{(X-X^{\prime})^{2}+(Y-Y^{\prime})^{2}}\big{)}+O(\log_{+}c)+O(1).$